OpenStudy (cookiemonster18):

2.Find the slope of the linear line between the points (10, -4) and (4, -4).

Mehek (mehek14):

\(\dfrac{y_2-y_1}{x_2-x_1}\) y2 = -4 y1 = -4 x2 = 4 x1 = 10

Mehek (mehek14):

\(\dfrac{-4-(-4)}{4-10}\)

OpenStudy (cookiemonster18):

so the answer is -6? because -4-(-4) equals 0 @Mehek14

Mehek (mehek14):

0/6 = 0

OpenStudy (pawanyadav):

If the line is parallel to x axis it's slope is 0 means y is not increasing or decreasing with x But if a line parallel to y-axis Then its slope is infinity.